OpenStudy (mollylasko):

If the margin of error in an estimate for the mean weight of a shipment is + or -2 pounds at a confidence level of 95 percent, what will be the margin of error at a confidence level of 98 percent? Be sure to show how you arrived at your answer.

jimthompson5910 (jim_thompson5910):

Recall that you'll use this table to find the critical z values https://www.ltcconline.net/greenl/courses/201/estimation/smallConfLevelTable.htm

jimthompson5910 (jim_thompson5910):

The margin of error M is equal to \[\Large M = z*\frac{s}{\sqrt{n}}\] The part \(\Large \frac{s}{\sqrt{n}}\) will stay the same. Only the value of z changes as the confidence level changes

OpenStudy (mollylasko):

i see that 1.96 is the confidence interval at 95%, so do you just put what is the interval at 98?

OpenStudy (mollylasko):

what is n?

jimthompson5910 (jim_thompson5910):

at 95% confidence, z = 1.96 s and n are unknown but that doesn't matter. All that matters is that they are fixed and don't change

jimthompson5910 (jim_thompson5910):

at 95% confidence, z = 1.96 and M = 2, so \[\Large M = z*\frac{s}{\sqrt{n}}\] \[\Large 2 = 1.96*\frac{s}{\sqrt{n}}\] agreed so far?

OpenStudy (mollylasko):

yes

OpenStudy (mollylasko):

so then you replace it with 2.33?

jimthompson5910 (jim_thompson5910):

ok let's isolate the \(\Large \frac{s}{\sqrt{n}}\) part

jimthompson5910 (jim_thompson5910):

\[\Large 2 = 1.96*\frac{s}{\sqrt{n}}\] \[\Large \frac{2}{1.96} = \frac{s}{\sqrt{n}}\] \[\Large 1.020408 = \frac{s}{\sqrt{n}}\] \[\Large \frac{s}{\sqrt{n}} = 1.020408\] agreed?

OpenStudy (mollylasko):

yes!

jimthompson5910 (jim_thompson5910):

so we don't know what s or n is, but we know that \[\Large \frac{s}{\sqrt{n}} = 1.020408\]

OpenStudy (mollylasko):

okay so then you multiply that with the 2.33?

jimthompson5910 (jim_thompson5910):

yes you'll multiply that with 2.33 \[\Large M = z*\frac{s}{\sqrt{n}}\] \[\Large M = 2.33*1.020408\] \[\Large M = ???\]

jimthompson5910 (jim_thompson5910):

this works because s & n are fixed, so the quantity \(\Large \frac{s}{\sqrt{n}}\) is also fixed

OpenStudy (mollylasko):

2.37766064?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so if you rounded to 2 decimal points, then it would be M = 2.38

OpenStudy (mollylasko):

yay!! thank you so much

jimthompson5910 (jim_thompson5910):

no problem